OpenStudy (anonymous):

How would I simplify 3√64(x+1)^3

OpenStudy (campbell_st):

well if you are looking at \[\sqrt[3]{64(x + 1)^3}\] then you need to know \[64 = 4^3\] so writing the problem in index form you have \[(4^3)^{\frac{1}{3}} \times((x + 1)^3)^{\frac{1}{3}}\] simplify by multiplying the powers...